TEX Users Group

P.O. Box 869 Santa Barbara, California 93102 USA

Board of Directors^{1}

Officers: Christina Thiele (President) Michel Goossens (Vice-President) George Greenwade (Treasurer) Peter Flynn (Secretary)

Jan 1993 – Jul 1995:  Jan 1994 – Jul 1997:
Christina Thiele^{2}  Luzia Dietsche

George Greenwade

Jan 1994 – Jul 1995:^{3}  Yannis Haralambous
Michel Goossens  Jon Radel
Michael Doob  Sebastian Rahtz

Tom Rokicki

Jan 1994 – Jul 1996:  Jul 1995–1998:^{4}
Mimi Burbank  Barbara Beeton^{5}
Jackie Damrau  Karl Berry
Michael Ferguson  Judy Johnson
Peter Flynn  Jiří Zlatuška
Nico Poppelier  [to be appointed]^{6}

Special Directors: Johannes Braams (NTG) Dag Langmyhr (Nordic Group) Chris Rowley (UKTUG)

1 After the 1993 fall election, board terms began in January 1994. Lots drawn to assign 5 members each to 1.5-, 2.5-, and 3.5-year terms, to achieve a 3-year term beginning/ending with the July board meetings (hence the references to both months in the list of terms of office).

2 Term of 2.5 years ends; new President’s term begins at July 1995 board meeting.

3 Appointed by the President, January 1994. There were only 11 candidates for the Fall 1993 election, so the remaining four slots were to be filled by one-year appointees. Only three appointments were made.

4 Terms to begin in July 1995 (results of Spring 1995 election); only four candidates presented themselves, therefore they win by acclamation, and a fifth must be named by the new President at the July 1995 board meeting.

5 Previous term: Jan 1994 – Jul 1995.

6 Appointed members serve only until the next scheduled election, which will take place in one year.

Tel.: +1 (805) 963-1338  tug@tug.org  Fax: +1 (805) 963-8358

Source: /tex-archive/usergrps/tug/tugboard.tex on CTAN. Rev. May 95. Next update: July 95.

Addresses of Board Members

Barbara Beeton:

bnb@math.ams.org

American Mathematical Society, P.O. Box 6248, Providence, RI 02940 USA, 401-455-4014

Karl Berry:

kb@cs.umb.edu

135 Center Hill Road, Plymouth, MA 02360 USA, 508-224-2367

Johannes Braams:

j.l.braams@research.ptt.nl

Kooienswater 62, NL-2715 AJ Zoetermeer, The Netherlands

Mimi Burbank:

mimi@scri.fsu.edu

Supercomputer Computations Research Institute (SCRI), B-186, 400 Science Center Library, Florida State University, Tallahassee, FL 32306-4052 USA, 904-644-2440

Jackie Damrau:

jdamrau@intervoice.com

P.O. Box 875, Red Oak, TX 75154-0875 USA, 214-617-2323

Luzia Dietsche:

luzia.dietsche@urz.uni-heidelberg.de

Universitätsrechenzentrum, Im Neuenheimer Feld 293, D-69120 Heidelberg, Germany

Michael Doob:

mdoob@cc.umanitoba.ca

Dept. of Math & Astronomy, 341 Machray Hall, University of Manitoba, Winnipeg, Manitoba, R3T 2N2, Canada, 204-474-9796

Michael Ferguson:

mike@inrs-telecom.uquebec.ca

INRS – Télécommunications, Université du Québec, 16 Place du Commerce, Verdun, Québec, H3E 1H6, Canada, 514-765-7834

Peter Flynn:

pflynn@curia.ucc.ie

Computer Center, University College, Cork, Ireland, +353 21 276871 x2609

Michel Goossens:

goossens@cern.ch

CN Division, CERN, CH-1211 Geneva 23, Switzerland, +41.22.767.3363, FAX +41.22.767.8630

George Greenwade:

bed gdg@shsu.edu

Dept. of Economics and Business Analysis, College of Business Administration, P.O. Box 2118, Sam Houston State University, Huntsville, TX 77341-2118 USA Yannis Haralambous:

                                            yannis.haralambous@univ-lille1.fr

187, rue Nationale, F-59000 Lille, France, FAX (33) 20.91.05.64

Judy Johnson:

                                            judy@etp.com

ETP Services, 2906 NE Glisan, Portland, OR 97232 USA

Dag Langmyhr:

                                            dag@ifi.uio.no

Dept. of Informatics, University of Oslo, P.O. Box 1080 Blindern, N-0316 Oslo, Norway, +47 22 85 24 50; FAX +47 22 85 24 01

Nico Poppelier:

                                            n.poppelier@elsevier.nl

Elsevier Science Publishers BV, Academic Publishing Division, I nformation Technology Development, Molenwerf 1, NL-1014 AG Amsterdam, The Netherlands

Jon Radel:

                                            jon@radel.com

P.O. Box 2276, Reston, VA 22090-0276 USA

Sebastian Rahtz:

                                            s.rahtz@elsevier.co.uk

7 Stratfield Road, Oxford OX2 7BG, UK, +44 1865 843662

Tom Rokicki:

                                            rokicki@cs.stanford.edu

Box 2081, Stanford, CA 94309 USA,

415-322-6442

Chris Rowley:

                                            c.a.rowley@open.ac.uk

Open University, Walton Hall, M ilton Keynes MK7 6AA, UK

Christina Thiele:

                                            cthiele@ccs.carleton.ca

15 Wiltshire Circle, Nepean, Ontario, K2J 4K9, Canada, 613-823-3630

Jiří Zlatuška:

                                            zlatuska@muni.cz

Faculty of Informatics, Masaryk University, Burešova 20, CZ-602 00 Brno, Czech Republic, +42-5-41213125 or +42-5-41213219, FAX +42-5-41212747